# Cycle Matroids of Graphings: From Convergence to Duality

### v2

@@ Abstract
In this paper, we study the connection between local-global convergence of graphs and quotient-convergence of their cycle matroids. We characterize the exposed points of the associated convex set, forming an analytic counterpart of matroid base-polytopes. Finally, we consider dual planar graphings and show that the cycle matroid of one is the cocycle matroid of its dual if and only if the underlying graphings are hyperfinite.
@@ 1 Introduction
Cycle matroids bridge graph theory and matroid theory by providing matroid theoretic counterparts of graph concepts such as cycles, cuts, and spanning trees, thus leading to a powerful framework for understanding and solving various combinatorial and optimization problems. The matroidal perspective allows for a deeper analysis of graph properties and leads to efficient algorithms in network design, electrical engineering, and operations research.
Graphings are bounded degree graphs on standard probability measure spaces, serving as limit objects in sparse graph limit theory, see Subsection 2.2 for the definition. Recently, Lovász [18] initiated the study of cycle matroids from an analytic point of view and introduced the cycle matroid of a graphing $\mathcal{G}=(X,E,\mu)$ by defining its rank function as
$\rho_{\mathcal{G}}(F)=1-\mathbb{E}_{x}\left[\frac{1}{|V(\mathcal{G}^{F}_{x})|}\right],$ (1)
@@
where $F\subseteq E$ is Borel, $x$ is a $\mu$-random vertex of $\mathcal{G}$ and $\mathcal{G}^{F}_{x}$ denotes its connected component in the graphing $\mathcal{G}^{F}=(X,F,\mu)$. It was shown that $\rho_{\mathcal{G}}$ is an increasing submodular setfunction on the Borel subsets of $E$. It is not difficult to verify that this definition generalizes the normalized rank of the cycle matroid of a finite graph.
@@
This raises several questions about cycle matroids of graphings. How to define and describe their bases, cycles, and other basic properties known from the finite case? Does the convergence of a sequence of finite graphs to a graphing imply convergence of the corresponding matroids (in the sense introduced in [3])? Does duality of matroids play any role? In this paper, we provide some answers to these questions. We develop the limit theory of these matroids, study their bases, and investigate how they behave under duality of planar graphings. We give some background on earlier work on cycle matroids of infinite graphs in Subsection 1.5.
1.1 Limit theory
@@
Our first result connects local-global convergence of graphs with quotient-convergence of the rank function of cycle matroids, defined in [3]. We give a short definition, but refer to Subsection 2.4 for the setup and a precise formulation. A $k$-quotient of a setfunction $\varphi$ on a set-algebra $(J,\mathcal{B})$ is $\varphi\circ F^{-1}$, where $F:J\to[k]$ is a measurable map. We denote by $Q_{k}(\varphi)$ the set of $k$-quotients of $\varphi$. A sequence of setfunctions $(\varphi_{1},\varphi_{2},\dots)$ quotient converges to $\varphi$, denoted $\varphi_{n}\rightarrowtail\varphi$, if $Q_{k}(\varphi)\to Q_{k}(\varphi)$ in the Hausdorff distance for all $k\in\mathbb{N}$.
@@
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to the graphing $\mathcal{G}$ in the local-global sense. Let $r_{G_{n}}$ denote the rank function of the cycle matroid of $G_{n}$. Then
$\frac{r_{G_{n}}}{|V(G_{n})|}\rightarrowtail\rho_{\mathcal{G}}$
@@
We also exhibit an example showing that assuming local convergence of $G_{n}$ is not enough.
@@
There exists a sequence $(H_{1},H_{2},\dots)$ of finite graphs with all degrees bounded by $D$ that converges to a graphing $\mathcal{G}$ in the local sense, however
[…]
1.2 Fractional bases
For a finite matroid $M=(E,r)$ with ground set $E$ and rank function $r$, the set of fractional bases
$C_{M}=\left\{x\in\mathbb{R}^{E}~{}\Big{|}~{}x\geq 0,\ \sum_{e\in A}x(a)\leq r(A)\text{ for }A\subseteq E,\ \sum_{e\in E}x(e)=r(E)\right\}$
forms an integral polytope whose vertices are exactly the bases of $M$, called the base-polytope of $M$. For the cycle matroid of a connected graph, this means that among fractional spanning trees, the extreme ones are genuine spanning trees.
In the analytic setting developed by Lovász [19], the analogue of a base-polytope is the set of basic minorizing finitely additive measures. In particular, for a graphing $\mathcal{G}=(X,E,\mu)$, this gives
$C_{\mathcal{G}}=\text{bmm}^{+}(\rho_{\mathcal{G}})=\big{\{}\alpha\textrm{ finitely additive measure on $E$}~{}\big{|}~{}\alpha\leq\rho_{\mathcal{G}},\ \alpha(E)=\rho_{\mathcal{G}}(E)\big{\}}.$
A vertex of a convex polytope generalizes to the infinite setting in several ways. A point $v$ of the convex set $C$ is extreme if it cannot be written as $v=\frac{1}{2}(u+w)$, where $u,w\in C\backslash\{v\}$. The point $v$ is exposed if there is a supporting hyperplane $H$ through $v$ such that $H\cap C=\{v\}$. This property depends on the topology of the ambient vector space; in this paper, we mean that $H$ is defined by the inner product with a function in $L^{1}(E)$. An exposed point is necessarily extreme, but not the other way around; see Section 4 for further comments.
An essential spanning forest of a graphing $\mathcal{G}$ is an acyclic subset $T$ of the edges that forms a spanning tree in finite components of $\mathcal{G}$, and contains only infinite $T$-components in the infinite $\mathcal{G}$-components. Note that $T$ might contain several components in an infinite component of $\mathcal{G}$, but all of those are required to be infinite; this is why $T$ is called a spanning forest instead of a spanning tree. As a generalization of the finite case, we prove the following.
@@ 1.2 Fractional bases
Let $\mathcal{G}$ be a graphing. Then, the exposed points of $C_{\mathcal{G}}$ are exactly the hyperfinite essential spanning forests of $\mathcal{G}$.
The following question remains open.
@@
Are all extreme points of $C_{\mathcal{G}}$ exposed, and therefore hyperfinite spanning forests?
1.3 Matroids beyond the hyperfinite
While the definition of the cycle matroid via (1) is a consistent generalization of the finite notion, it does have a drawback: the rank of the full edge set is at most 1. This is natural for finite graphs, as spanning trees have average degree slightly below 2. For graphings, however, spanning trees can have higher average degree. The notion of cost is designed to measure the minimal amount of edges needed to generate the connected components of a graphing; see the precise definition in Subsection 2.5. It seems natural to ask if one can come up with a notion of cycle matroid whose total rank is equal to the cost.
Question 1.5.
Is there a notion of the cycle matroid of a graphing $\mathcal{G}$ with submodular rank function $\varphi_{\mathcal{G}}$ such that $\varphi_{\mathcal{G}}(E)=\textrm{cost}(\mathcal{G})$?
A positive answer to this question would not invalidate $\rho_{\mathcal{G}}$, rather complement it. See e.g. Theorems 1.9 and 1.10 below, where we use ideas from [7] to give a positive answer for a certain class of planar graphings.
To further justify the study of $\rho_{\mathcal{G}}$, we highlight that even though $\rho_{\mathcal{G}}$ only sees spanning forests that are hyperfinite, it still carries a lot of information about the graphing. In the finite case, Whitney proved that every 3-connected graph is uniquely determined by its cycle matroid. In an upcoming work [2], we study analogous problems for graphings.
Question 1.6.
Under what assumptions does $\rho_{\mathcal{G}}$ determine the graphing $\mathcal{G}$ (up to measure zero)?
As a first step in this direction, we could show that $\rho_{\mathcal{G}}$ already determines acyclic graphings with minimum degree 3 up to isomorphism in the following sense: For a graphing $\mathcal{G}$ and an acyclic graphing $\mathcal{T}$ of minimum degree 3, if $\rho_{\mathcal{G}}\cong\rho_{\mathcal{T}}$ then $\mathcal{G}\cong\mathcal{T}$. In both cases, $\cong$ means that there exists a measure preserving bijection between the ground spaces, up to measure zero, that maps one structure to the other.
1.4 Duality
Duality is an important concept of matroid theory, and it lines up nicely with duality of finite planar graphs: the dual of a cycle matroid of an (embedded) planar graph is the cycle matroid of its planar dual. The question naturally arises: Does an analogous statement hold for planar graphings?
The notion of a planar graphing takes some work to set up. We postpone this to Section 5, where we give an informal introduction and refer to [1] and [7] for a rigorous treatment. For now, we only say that in a planar graphing, the connected components are planar maps, i.e. they are endowed with a specified planar embedding. We assume that these planar maps have no accumulation points in the plane. We also want to point out that the notion of a “dual graph” is not completely evident when infinite faces are present [24, 25].
When we talk about a pair of dual graphings $\mathcal{G}$ and $\mathcal{G}^{*}$, it is assumed, informally speaking, that the components of $\mathcal{G}^{*}$ are, as planar maps, the duals of the components of $\mathcal{G}$. For simplicity, we exclude infinite faces, we assume that components are $2$-vertex-connected to avoid edges that are only adjacent to one face, and that the maximum size of the faces is bounded to make sure that both $\mathcal{G}$ and $\mathcal{G}^{*}$ have bounded degree. For a dual pair, there is a Borel bijection $\sigma:~{}E(\mathcal{G})\to E(\mathcal{G}^{*})$ such that the image of the star of a vertex of $\mathcal{G}$ is a facial cycle in $\mathcal{G}^{*}$ and vice versa. We also note that the hyperfiniteness of $\mathcal{G}$ and $\mathcal{G}^{*}$ are equivalent [1, Example 3.11, Proposition 3.12].
@@ 1.4 Duality
The dual of the cycle matroid, called the cocyclic matroid of $\mathcal{G}$, can be defined for any (not necessarily planar) graphing by its rank function
$\rho^{D}_{\mathcal{G}}(A)=\rho_{\mathcal{G}}(E\setminus A)+\tilde{\mu}(A)-\rho_{\mathcal{G}}(E),$
where $\tilde{\mu}$ denotes the edge measure of $\mathcal{G}$. Our first result here shows that in the hyperfinite case duality works as in the finite case.
@@
Let $\mathcal{G}$ and $\mathcal{G}^{*}$ be a hyperfinite pair of dual planar graphings. Then, $\rho^{D}_{\mathcal{G}}=\rho_{\mathcal{G}^{*}}$ and $\rho^{D}_{\mathcal{G}^{*}}=\rho_{\mathcal{G}}$.
However, when $\mathcal{G}$ and $\mathcal{G}^{*}$ are not hyperfinite, the two matroids do not coincide. This is the case for example if the components of $\mathcal{G}$ and $\mathcal{G}^{*}$ are dual hyperbolic tilings. We can actually exploit this phenomenon to give a partial positive answer to Question 1.5.
@@
Let $\mathcal{G}$ and $\mathcal{G}^{*}$ be a non-hyperfinite pair of dual planar graphings. Then, $\rho^{D}_{\mathcal{G}^{*}}\neq\rho_{\mathcal{G}}$. Moreover, $\rho^{D}_{\mathcal{G}^{*}}(E)$ is a submodular function on $E(\mathcal{G})$ that attains the cost, i.e.
$\rho^{D}_{\mathcal{G}^{*}}(E)=\textrm{cost}(\mathcal{G}).$
Note that planar graphings are treeable [7], so if they are not hyperfinite, their cost is strictly more than $1$. Both submodular functions $\rho_{\mathcal{G}}$ and $\rho^{D}_{\mathcal{G}^{*}}$ are natural to consider, as supported by the following result.
@@
Let $\mathcal{G}$ and $\mathcal{G}^{*}$ be a pair of dual planar graphings. Let $<$ be an arbitrary Borel total order on the edges. The Wired Maximal Spanning Forest $\mathcal{T}^{<}_{w}$ is an exposed point of $\text{bmm}^{+}(\rho_{\mathcal{G}})$, while the Free Maximal Spanning Forest $\mathcal{T}^{<}_{f}$ is an exposed point of $\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$.
In [1, Theorem 1 (15)], it was shown that a unimodular random map is hyperfinite if and only if the wired and free maximal spanning forests coincide. In our setup, this can be interpreted as $\mathcal{T}^{<}_{f}=\mathcal{T}^{<}_{w}$ in the special case when $\mathcal{G}$ is an (edge) Bernoulli graphing over the unimodular random planar map, and $<$ is the natural ordering on the edges given by the edge-labeling.
1.5 Other notions of cycle matroids on infinite graphs
In this paper we study the cycle matroids of graphings, as graphings are natural limit objects in the context of sparse graph limits. It is equally natural to extend the notion of cycle matroid to infinite graphs without a measurable structure, for example connected locally finite graphs. There has been an extensive line of research on this topic. Although it is not directly related to our results, we mention some of the highlights.
Probably the simplest notion for infinite graphs is the finitary cycle matroid whose circuits are exactly the finite cycles. However, an unfortunate feature of finitary matroids is that excluding infinite circuits spoils duality. As a workaround, Bruhn, Diestel, Kriesell, Pendavingh and Wollan [6] proposed five equivalent sets of axioms leading to a non-finitary theory of infinite matroids. These axioms were in fact shown to be equivalent to the notion of B-matroids introduced by Higgs [11]. Using these axioms, Bruhn and Diestel [5] provided further examples by defining the circuits of the matroid to be the edge sets of topological circuits of compactifications of infinite graphs obtained by adding the space of ends to the graph. Later, Bowler, Carmesin and Christian [4] introduced a class of infinite cycle matroids that unifies previous constructions.

### v3

@@ Abstract
In this paper, we study the connection between local-global convergence of graphs and quotient-convergence of their cycle matroids. We characterize the exposed points of associated convex sets, forming an analytic counterpart of matroid independence- and base-polytopes. Finally, we consider dual planar graphings and show that the cycle matroid of one is the cocycle matroid of its dual if and only if the underlying graphings are hyperfinite.
@@ 1 Introduction
Cycle matroids bridge graph theory and matroid theory by providing matroid theoretic counterparts of graph concepts such as cycles, cuts, and spanning trees, thus leading to a powerful framework for understanding and solving various combinatorial and optimization problems. The matroidal perspective allows for a deeper analysis of graph properties and leads to efficient algorithms in network design, electrical engineering, and operations research. In particular, the cycle matroid of a graph $G$ is characterized by its rank function, where the rank of a set $F\subseteq E(G)$ of edges is $r_{G}(F)=|V(G)|-|\{\textrm{$F$-connected components of $G$}\}|$; see Section 2 for properties of matroid rank functions.
There has been an extensive line of research on extending the notion of cycle matroid to infinite graphs, for example for connected locally finite graphs. Probably the simplest notion for infinite graphs is the finitary cycle matroid whose circuits are exactly the finite cycles. However, an unfortunate feature of finitary matroids is that excluding infinite circuits spoils duality. As a workaround, Bruhn, Diestel, Kriesell, Pendavingh and Wollan [9] proposed five equivalent sets of axioms leading to a non-finitary theory of infinite matroids. These axioms were in fact shown to be equivalent to the notion of B-matroids introduced by Higgs [14]. Using these axioms, Bruhn and Diestel [8] provided further examples by defining the circuits of the matroid to be the edge sets of topological circuits of compactifications of infinite graphs obtained by adding the space of ends to the graph. Later, Bowler, Carmesin and Christian [7] introduced a class of infinite cycle matroids that unifies previous constructions. Motivated by the work of Nešetřil and Ossona de Mendez [26] on first order convergence of graphs, Kardoš, Král’, Liebenau, and Mach [15] introduced the notion of first order convergence of linearly representable matroids. Since graphic matroids are linearly representable, this also provides a convergence notion for such matroids.
In addition to the results discussed above, however, there have been no or only moderate work concerning graphs with an underlying measurable structure. In this paper we study the cycle matroids of graphings, as graphings are natural limit objects in the context of sparse graph limits. A graphing is a bounded degree graph $\mathbf{G}=(J,\mu,E)$, where $(J,\mu)$ is a standard Borel probability space, the edge set $E$ is a Borel subset of $\binom{J}{2}$ with a certain measure preserving property; this property allows one to define the measure $\eta_{\mathbf{G}}$ on the edge set, see Subsection 2.2 for the full definition. Graphings serve as limit objects in sparse graph limit theory. Recently, Lovász [23] initiated a limit theory of cycle matroids of bounded-degree graphs, and introduced the cycle matroid of a graphing $\mathbf{G}=(J,\mu,E)$ by defining its rank function as
$\rho_{\mathbf{G}}(F)=1-\mathbb{E}_{x}\left[\frac{1}{|V(\mathbf{G}^{F}_{x})|}\right],$ (1)
@@
where $F\subseteq E$ is Borel, $x$ is a uniform random point from $\mu$, $\mathbf{G}^{F}_{x}$ denotes its connected component in the graphing $\mathbf{G}^{Y}=(J,\mu,F)$, and $V(\mathbf{G}^{F}_{x})$ denotes the set of vertices of the component. In [23], it was shown that $\rho_{\mathbf{G}}$ is an increasing submodular setfunction on the Borel subsets of $E$, and $\rho_{\mathbf{G}}\leq\eta_{\mathbf{G}}$. It is not difficult to verify that this definition generalizes the rank function of the cycle matroid of a finite graph, normalized by the number of vertices. In particular, if $\mathbf{G}$ a forest with finite components, then $\rho_{\mathbf{G}}(E)$ is half of the average degree. If $\mathbf{G}$ is any graphing with infinite components, then $\rho_{\mathbf{G}}(E)=1$.
@@
This raises several questions about cycle matroids of graphings. How to define and describe their independent sets, bases, and other basic properties known from the finite case? Does the convergence of a sequence of finite graphs to a graphing imply convergence of the corresponding matroids (in the sense introduced in [5])? Does duality of matroids play any role? In this paper, we provide some answers to these questions. We develop the limit theory of cycle matroids of bounded-degree graphs, show that cycle matroids of graphings are their limit objects in the sense of [5], we study their bases, and investigate how they behave under duality of planar graphings.
Limit theory.
@@
Our first result connects local-global convergence of graphs with quotient-convergence of the rank function of cycle matroids, defined in [5]. A $k$-quotient of a setfunction $\varphi$ on a set-algebra $(J,\mathcal{B})$ is $\varphi\circ F^{-1}$, where $F:J\to[k]$ is a measurable map. We denote by $Q_{k}(\varphi)$ the set of $k$-quotients of $\varphi$. A sequence of setfunctions $(\varphi_{1},\varphi_{2},\dots)$ quotient converges to $\varphi$, denoted $\varphi_{n}\rightarrowtail\varphi$, if $Q_{k}(\varphi_{n})\to Q_{k}(\varphi)$ as $n\to\infty$ in the Hausdorff distance for all $k\in\mathbb{N}$. Note that $Q_{k}(\varphi_{n})$ and $Q_{k}(\varphi)$ are subsets of the same $2^{k}$-dimensional linear space, so convergence in Hausdorff distance does not depend on the norm of this space. We refer to Subsection 2.4 for more details on this convergence notion. Our first main theorem concerns the case of cycle matroids of bounded-degree graphs.
@@
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to the graphing $\mathbf{G}$ in the local-global sense. Let $r_{G_{n}}$ denote the rank function of the cycle matroid of $G_{n}$. Then
$\frac{r_{G_{n}}}{|V(G_{n})|}\rightarrowtail\rho_{\mathbf{G}}$
@@
[…]
Independent sets.
Our goal is to define independent sets of the cycle matroid of a graphing. There are various ways to characterize independent sets in the cycle matroid of a finite graph, of which we consider the following four.
  • Independent sets in the cycle matroid of a finite graph are edge-sets of acyclic subgraphs (briefly, forests).
  • Bases of the cycle matroid of a finite graph are spanning forests, i.e. they are spanning trees in every connected component. The independent sets are their subsets.
  • Independent sets in any matroid can be recovered from the rank function as sets with $r(X)=|X|$.
  • In a finite matroid $(E,r)$, solutions of the system of inequalities $x_{e}\geq 0$ $(e\in E)$, and $\sum_{e\in A}x(a)\leq r(A)$ $(A\subseteq E)$ are called fractional independent sets, and form the matroid polytope. Vertices of the matroid polytope are precisely the (indicator vectors of) independent sets.
Before generalizing these definitions for graphings and stating the theorem about a certain equivalence of them, we need to discuss some points where the generalizations are not quite straightforward.
In this section, we consider forests that are hyperfinite. A graphing $\mathbf{G}=(J,\mu,E)$ is hyperfinite if, for every $\varepsilon>0$, there is a Borel set $S\subseteq E$ with $\eta_{\mathbf{G}}(S)\leq\varepsilon$ such that every connected component of $E\setminus S$ is finite. A forest in $\mathbf{G}$ is an acyclic subset $T\subseteq E$ of the edges. An essential spanning forest is a forest that forms a spanning tree in every finite component of $\mathbf{G}$, and contains only infinite $T$-components in the infinite $\mathbf{G}$-components. Note that $T$ might contain several components in an infinite component of $\mathbf{G}$, but all of those are required to be infinite; this is why $T$ is called an essential spanning forest instead of a spanning tree. Let us emphasize that we will call a set of edges a “forest” or an “essential spanning forest” if they became one after adding or deleting a set of edges of measure 00.
A vertex of a convex polytope can be generalized to the infinite setting in several ways. A point $v$ of the convex set $C$ in a Banach space $X$ is extreme, if it cannot be written as $v=\frac{1}{2}(u+w)$, where $u,w\in C\backslash\{v\}$. The point $v$ is exposed, if there is a continuous linear functional $\Lambda$ on $X$ such that $\Lambda(y)\leq\Lambda(x)$ for every $y\in C$, and $x$ is the unique point of $C$ with equality. An exposed point is necessarily extreme, but not the other way around; see Section 4 for further comments.
It was suggested in [22] that for a general increasing submodular setfunction $\varphi$ with $\varphi(\emptyset)=0$, the generalization of the matroid polytope should be the set of its minorizing charges, i.e. finitely additive measures $\alpha$ such that $0\leq\alpha\leq\varphi$. For a graphing $\mathbf{G}=(J,\mu,E)$, we are in a simpler situation, since a finitely additive measure $\alpha$ minorizing $\rho_{\mathbf{G}}$ satisfies $\alpha\leq\eta_{\mathbf{G}}$, and so it is automatically countably additive (i.e., a proper measure). So the analytic version of the cycle matroid polytope will be
$\text{\sf mm}_{+}(\rho_{\mathbf{G}})=\big{\{}\alpha\textrm{ measure on $E$}\mid 0\leq\alpha\leq\rho_{\mathbf{G}}\big{\}}.$
The base polytope of a cycle matroid can be generalized as
$\text{\sf bmm}_{+}(\rho_{\mathbf{G}})=\big{\{}\alpha\in\text{\sf mm}_{+}(\rho_{\mathbf{G}})\mid\alpha(E)=\rho_{\mathbf{G}}(E)\big{\}}.$
First, we give a characterization of the exposed points of $\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$.
@@
Let $\mathbf{G}=(J,\mu,E)$ be a graphing.
  1. (a)
    $\alpha\in\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ is exposed if and only if $\alpha=\eta_{\mathbf{G}}|_{F}$ for a hyperfinite subforest $F\subseteq E$.
  2. (b)
    $\alpha\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ is exposed if and only if $\alpha=\eta_{\mathbf{G}}|_{F}$ for a hyperfinite essential spanning forest $F\subseteq E$.
We also give characterization of these kind of subsets used in Theorem 1.2.
@@
Let $\mathbf{G}=(J,\mu,E)$ be a graphing and $F\subseteq E$ be a Borel subset of edges. Then the following are equivalent:
  1. (i)
    $(J,\mu,F)$ is a hyperfinite forest;
  2. (ii)
    $F$ is a Borel subset of the edge set of a hyperfinite essential spanning forest;
  3. (iii)
    $\rho_{\mathbf{G}}(F)=\eta_{\mathbf{G}}(F)$.
Similarly, the following are equivalent:
  1. (i’)
    $(J,\mu,F)\subseteq(J,\mu,E)$ is a hyperfinite essential spanning forest;
  2. (ii’)
    $\rho_{\mathbf{G}}(F)=\eta_{\mathbf{G}}(F)=\rho_{\mathbf{G}}(E)$.
Along the lines, we introduce the notion of weakly exposed points and prove structural properties of those, see Theorem 4.7 and Corollary 4.8. In Section 4.5, we show that hyperfinite forests satisfy an exchange property, which in turn leads to a conceptually new proof of the submodularity of the set function $\rho_{\mathbf{G}}$, different from the (somewhat lengthy) proof in [23]. The following question remains open.
@@
Are all extreme points of $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ exposed, and therefore hyperfinite spanning forests?
Duality.
Duality is an important concept of matroid theory, and it lines up nicely with duality of finite planar graphs: the dual of the cycle matroid of a planar map (embedded connected planar graph) is the cycle matroid of its planar dual. The question naturally arises: Does an analogous statement hold for planar graphings?
The notion of a “planar graphing” takes some work to set up. In order to talk about pairs of dual planar graphings, we define planar graphings $\mathbf{G}=(J,\mu,E)$ in a rather restricted sense. The connected components of a graphing map are 2-vertex-connected graphs endowed with a specified planar embedding. We assume that the vertex sets of these planar maps have no accumulation points in the plane, and all faces are finite cycles, whose lengths are uniformly bounded. Then the dual graphing map $\mathbf{G}^{*}=(J^{*},\mu^{*},E^{*})$ is well defined, and it has bounded degrees. For a dual pair of graphing maps, there is a measure preserving Borel bijection $\sigma\colon E\to E^{*}$ such that the image of the star of a vertex of $\mathbf{G}$ is a facial cycle in $\mathbf{G}^{*}$ and vice versa. We also note that the hyperfiniteness of $\mathbf{G}$ and of $\mathbf{G}^{*}$ are equivalent [3, Example 3.11, Proposition 3.12]. See Section 5 for an exact definition, and also [3] and [10] for a more general treatment.
@@
The dual of the cycle matroid, called the cocycle matroid of $\mathbf{G}=(J,\mu,E)$, can be defined for any (not necessarily planar) graphing by its rank function
$\rho^{*}_{\mathbf{G}}(A)=\rho_{\mathbf{G}}(E\setminus A)+\eta_{\mathbf{G}}(A)-\rho_{\mathbf{G}}(E).$
For a general notion of duality for submodular setfunctions, see [6]. Our first result here shows that in the hyperfinite case duality works similarly as in the finite case.
@@
Let $\mathbf{G}$ and $\mathbf{G}^{*}$ be a hyperfinite pair of dual planar graphings. Then $\rho^{*}_{\mathbf{G}}=\rho_{\mathbf{G}^{*}}$ and $\rho^{*}_{\mathbf{G}^{*}}=\rho_{\mathbf{G}}$.
However, when $\mathbf{G}$ and $\mathbf{G}^{*}$ are not hyperfinite, the two matroids do not form a dual pair. This is the case, for example, if the components of $\mathbf{G}$ and $\mathbf{G}^{*}$ are dual hyperbolic tilings. This example leads us to an interesting area, motivated by group theory, namely the theory of “cost”. The cost of a graphing $\textrm{\rm cost}(\mathbf{G})$ is a nonnegative number, designed to measure the minimal amount of edges needed to generate the connected components of a graphing; see the precise definition in Subsection 2.2. Every hyperfinite graphing has cost at most $1$, and exactly $1$ if all components are infinite. On the other hand, if a planar graphing is not hyperfinite, then its cost is strictly more than $1$ [10].
@@
Let $\mathbf{G}$ and $\mathbf{G}^{*}$ be a non-hyperfinite pair of dual planar graphings. Then $\rho^{*}_{\mathbf{G}^{*}}\neq\rho_{\mathbf{G}}$. Moreover, $\rho^{*}_{\mathbf{G}^{*}}$ is a submodular function on $E$ that attains the cost, i.e. $\rho^{*}_{\mathbf{G}^{*}}(E)=\textrm{\rm cost}(\mathbf{G})$.
Both submodular functions $\rho_{\mathbf{G}}$ and $\rho^{*}_{\mathbf{G}^{*}}$ are natural to consider, as supported by the following. Let $\mathbf{G}$ be a graphing and $<$ be an arbitrary Borel total order on the edges. We define the free maximal spanning forest $\mathbf{T}_{f}$ of $\mathbf{G}$ by simultaneously deleting all edges $e$ of $\mathbf{G}$ for which there exists a cycle $C$ containing $e$ in which $e$ is $<$-minimal. The wired maximal spanning forest $\mathbf{T}_{w}$ is similar to the free version, but we also remove any edge that is minimal with respect to the ordering on a two-way infinite path. It is straightforward to check that both $\mathbf{T}_{f},\mathbf{T}_{w}$ are acyclic and $\mathbf{T}_{w}\subseteq\mathbf{T}_{f}$. We refer to [24, Section 11] for a thorough introduction. It is also true that $\mathbf{T}_{w}$ is an essential spanning forest of $\mathbf{G}$. We show the following.
@@
Let $\mathbf{G}$ and $\mathbf{G}^{*}$ be a pair of dual planar graphings and $<$ be an arbitrary Borel total order on the edges. The Wired Maximal Spanning Forest $\mathbf{T}^{<}_{w}$ is an exposed point of $\text{\sf bmm}^{+}(\rho_{\mathbf{G}})$, while the Free Maximal Spanning Forest $\mathbf{T}^{<}_{f}$ is an exposed point of $\text{\sf bmm}^{+}(\rho^{*}_{\mathbf{G}^{*}})$.
Matroids beyond the hyperfinite.
While the definition of the cycle matroid via (1) is a consistent generalization of the finite notion, it does have a drawback: the rank of the full edge set is at most $1$. This is natural for finite graphs, as spanning trees have average degree slightly below 2. For graphings, however, spanning trees can have higher average degree, so if there is a notion of a matroid whose bases are spanning trees instead of essential spanning trees, then the total rank should be larger than $1$. Theorems 1.7 and 1.8 above show that for planar graphings, the dual matroid of the dual graphing provides such a submodular rank function.
To further justify the study of $\rho_{\mathbf{G}}$, we highlight that even though $\rho_{\mathbf{G}}$ only sees spanning forests that are hyperfinite, it still carries a lot of information about the graphing. In the finite case, Whitney proved that every 3-connected graph is uniquely determined by its cycle matroid. In an upcoming work [4], we study analogous problems for graphings.
Question 1.9.
Under what assumptions does $\rho_{\mathbf{G}}$ determine the graphing $\mathbf{G}$?
As a first step in this direction, we could show that $\rho_{\mathbf{G}}$ already determines acyclic graphings with minimum degree 3 up to isomorphism in the following sense: For a graphing $\mathbf{G}$ and an acyclic graphing $\mathbf{T}$ of minimum degree 3, if $\rho_{\mathbf{G}}\cong\rho_{\mathbf{T}}$ then $\mathbf{G}\cong\mathbf{T}$. In both cases, $\cong$ means that there exists a measure preserving bijection between the ground spaces, up to measure zero, that maps one structure to the other.